how to find the perimeter of a rectangle given area as 80 meters square and length as 16 meters
Answer by John10(297)

how to find the perimeter of a rectangle given area as 80 meters square and length as 16 meters
________________________________________________
Hello:)
You know area of rectangle is length times width
A = l * w = 80
And the length is 16: l = 16
then the width is: w = area/length = 80/16 = 5
The perimeter is P = 2(l + w)
You can finish from there :)
John10